# Fixed Points and Cycles of Parking Functions

### v1

@@ 3. Prime parking functions
As with classical parking functions, we can also study prime parking functions via circular rotation. The modified circular symmetry argument was first due to Kalikow [7, pp. 141-142], who provided the following observation: Let $G$ denote the group of all $n$-tuples $(a_{1},\dots,a_{n})\in[n-1]^{n}$ with componentwise addition modulo $n-1$. Let $H$ be the subgroup generated by $(1,1,\dots,1)$. Then every coset of $H$ contains exactly one prime parking function. The results we present in this section are thus largely parallel to the results for classical parking functions in our previous section.
@@
$\sum_{\ell=0}^{n-k-1}(-1)^{\ell}\binom{n-1}{k,\ell,n-k-\ell-1}(n-1)^{n-k-\ell-1}=\binom{n-1}{k}(n-2)^{n-k-1}.$
@@ 4. Further discussions
Fix some positive integers $k$ and $r$. Define an $(r,k)$-parking function of length $n$ [8] to be a sequence $(\pi_{1},\dots,\pi_{n})$ of positive integers such that if $\lambda_{1}\leq\cdots\leq\lambda_{n}$ is the (weakly) increasing rearrangement of $\pi_{1},\dots,\pi_{n}$, then $\lambda_{i}\leq k+(i-1)r$ for $1\leq i\leq n$. Denote the set of $(r,k)$-parking functions of length $n$ by $\operatorname{PF}_{n}(r,k)$. There is a similar interpretation for such $(r,k)$-parking functions in terms of the classical parking function scenario: One wishes to park $n$ cars on a street with $k+(n-1)r$ spots, but only $n$ spots are still empty, which are at positions no later than $k,\dots,k+(n-1)r$.
@@
We apply the generalized circle argument as described above. Arrange $k+nr$ spots in a circle. We first select $s$ spots for the first $s$ cars, which can be done in $\binom{k+nr}{s}$ ways. Then for the remaining $n-s$ cars, there are $(k+nr)^{n-s}$ possible preference sequences. Out of the $k+nr$ rotations for any preference sequence, exactly $k$ rotations become valid $(r,k)$-parking functions. The rest is standard circular symmetry argument. ∎
@@ acknowledgements
We thank Richard Stanley for helpful discussions. Some of the results in this paper were independently conjectured by Ari Cruz.

### v3

@@ 3. Prime parking functions
As with classical parking functions, we can also study prime parking functions via circular rotation. The modified circular symmetry argument was first due to Kalikow [15, pp. 141-142], who provided the following observation: Let $G$ denote the group of all $n$-tuples $(a_{1},\dots,a_{n})\in[n-1]^{n}$ with componentwise addition modulo $n-1$. Let $H$ be the subgroup generated by $(1,1,\dots,1)$. Then every coset of $H$ contains exactly one prime parking function. The results we present in this section are thus largely parallel to the results for classical parking functions in our previous section.
@@
$\sum_{\ell=0}^{n-k-1}(-1)^{\ell}\binom{n-1}{k,\ell,n-k-\ell-1}(n-1)^{n-k-\ell-1}$
$=$ $\binom{n-1}{k}\sum_{\ell=0}^{n-k-1}(-1)^{\ell}\binom{n-k-1}{\ell}(n-1)^{n-k-1-\ell}$
$=$ $\binom{n-1}{k}(n-2)^{n-k-1}.\hskip 56.9055pt\qed$
Theorem 3.2.
The generating function for prime parking functions of length $n$ with respect to the number of fixed points is given by $(q+n-2)^{n-1}$.
Proof.
Utilizing Theorem 3.1, we perform generating function calculations.
$\sum_{k=0}^{n-1}q^{k}\binom{n-1}{k}(n-2)^{n-k-1}=(q+n-2)^{n-1}.$
@@
Corollary 3.6.
Let $\pi\in\operatorname{PPF}_{n}$ be a prime parking function chosen uniformly at random. The expected number of cyclic points is given by
$\sum_{m=1}^{n-1}m!\binom{n-1}{m}/(n-1)^{m}.$
Proof.
As in the proof of Corollary 2.6, we apply Theorem 3.4 and note that every $m$-cycle contains $m$ cyclic points. ∎
@@ 4. Further discussions
4.1. Extensions of our techniques
Fix some positive integers $k$ and $r$. Define an $(r,k)$-parking function of length $n$ [16] to be a sequence $(\pi_{1},\dots,\pi_{n})$ of positive integers such that if $\lambda_{1}\leq\cdots\leq\lambda_{n}$ is the (weakly) increasing rearrangement of $\pi_{1},\dots,\pi_{n}$, then $\lambda_{i}\leq k+(i-1)r$ for $1\leq i\leq n$. Denote the set of $(r,k)$-parking functions of length $n$ by $\operatorname{PF}_{n}(r,k)$. There is a similar interpretation for such $(r,k)$-parking functions in terms of the classical parking function scenario: One wishes to park $n$ cars on a street with $k+(n-1)r$ spots, but only $n$ spots are still empty, which are at positions no later than $k,k+r,\dots,k+(n-1)r$.
@@ 4.1. Extensions of our techniques
We apply the generalized circle argument as described above. Arrange $k+nr$ spots in a circle. We first select $s$ spots for the first $s$ cars, which can be done in $\binom{k+nr}{s}$ ways. Then for the remaining $n-s$ cars, there are $(k+nr)^{n-s}$ possible preference sequences. Out of the $k+nr$ rotations for any preference sequence, exactly $k$ rotations become valid $(r,k)$-parking functions. The rest is a standard circular symmetry argument as in the proof of Theorems 2.1 and 2.4. ∎
4.2. Connections to other research areas
By setting $k=0,\dots,n$ in Theorem 2.1, a plethora of findings emerge, particularly through OEIS entries.
  • For $k=0$, our sequence is included as [FindStat, St001903] with explanation and coincides with [OEIS, A081215].
  • For $k=1$, our sequence coincides with [OEIS, A081216] and is related to the dimension of the primitive middle cohomology of Dwork hypersurfaces.
  • For $k=n-2$, our sequence coincides with [OEIS, A006325] which is the $4$-dimensional analog of centered polygonal numbers.
  • For $k=n-1$, our sequence coincides with [OEIS, A000217] which counts many interesting combinatorial objects, including the number of edges in a complete graph of order $n+1$ as well as the number of legal ways to insert a pair of parentheses in a string of $n$ letters.
[…]
We have not found a match on OEIS for other more general $k$ values, but the above list already presents many intriguing research directions, which we hope to explore in future work.
Likewise, setting $k=0,\dots,n-1$ in Theorem 3.1 reveals mysterious coincidences that motivate further investigations.
  • For $k=0$, our sequence coincides with [OEIS, A007778] which is the number of ways of writing an $n$-cycle as the product of $n+1$ transpositions.
  • For $k=1$, our sequence coincides with [OEIS, A055897] which is the total number of leaves in all labeled rooted trees with $n$ nodes.
  • For $k=2$, our sequence coincides with [OEIS, A081132] which is the sum of all the fixed points in the set of endofunctions on $\{1,2,\dots,n+1\}$.
  • For $k=n-3$, our sequence coincides with [OEIS, A019582] which counts many interesting combinatorial objects, including half the number of colorings of $4$ points on a line with $n$ colors as well as the number of ways to place two dominoes horizontally in different rows on an $n\times n$ chessboard.
  • For $k=n-2$, our sequence coincides with [OEIS, A002378] which are commonly referred to as the oblong numbers.
[…]
@@ Acknowledgements
We thank Richard Stanley for the helpful discussions. Some of the results in this paper were independently conjectured by Ari Cruz. The authors are grateful for many valuable comments from the referees, which significantly improved the quality of the paper.